OpenStudy (anonymous):

Identify the axis of symmetry for the quadratic function. y = x^2 + 2x – 3

OpenStudy (mertsj):

The x coordinate of the vertex is x = -b/2a. Since the axis of symmetry is a vertical line that passes through the vertex, that is also the equation of the axis of symmetry.

OpenStudy (anonymous):

So it would be -2/2(1) , then -2/2 ?

OpenStudy (mertsj):

So it would be x = -2/2 x=-1

OpenStudy (nathan917):

some how i got (1,-2) let me redo mine..

OpenStudy (anonymous):

but how did it become negative ?

OpenStudy (nathan917):

The graph would look like this?

OpenStudy (mertsj):

b=2, therefore -b = -2

OpenStudy (anonymous):

Oooooh okay I see now

OpenStudy (anonymous):

Thank You!(: I realize how to do it now!

OpenStudy (mertsj):

yw